OpenStudy (anonymous):

Use the properties of logarithms to condense the logarithmic expression. Write the expression as a single logaritm whose coefficient is 1. where possible, evaluate logarithmic expressions. logx+log(x^2-81)-log8-log(x+9) = ?

OpenStudy (campbell_st):

rewrite the expression log(x) + log(x^2 -81) - (log(8) + log(x +9) then using log law for addition log(x(x^2 -81) -(log(8(x +9)) which can be written as log(x(x-9)(x+9)) - log(8(x +9) now use the log law for division \[\log(\frac{x(x-9)(x+9)}{8(x + 9)})\] remove the common factor for the answer

OpenStudy (anonymous):

log ( x(x-9) / 8) ?

OpenStudy (anonymous):

YES..

OpenStudy (anonymous):

why thank you

OpenStudy (anonymous):

last question ... Evaluate or simplofy the expression without using a calculator ln e^3x

OpenStudy (anonymous):

\[\large \ln(e)^{3x} \implies 3x(\ln(e)) \implies 3x\]

OpenStudy (anonymous):

okay thanks a lot guys

OpenStudy (anonymous):

Welcome dear..

OpenStudy (anonymous):

im cramming so much in for a test